### v1

# Interim report for the International Muon Collider Collaboration

Abstract
This document summarises the International Muon Collider Collaboration (IMCC) progress and status of the Muon Collider R&D programme.
keywords:
Accelerator physics, Muon colliders, Accelerator and Detector Technology, Higgs Physics, Physics beyond the Standard Model (BSM).
Interim report for the
International Muon Collider Collaboration (IMCC)
C. Accettura1, S. Adrian2, R. Agarwal3, C. Ahdida1, C. Aimé4, A. Aksoy1,5, G. L. Alberghi6, S. Alden7, N. Amapane8,9, D. Amorim1, P. Andreetto10, F. Anulli11, R. Appleby12, A. Apresyan13, P. Asadi14, M. Attia Mahmoud15, B. Auchmann16,1, J. Back17, A. Badea18, K. J. Bae19, E. J. Bahng20, L. Balconi21,22, F. Balli23, L. Bandiera24, C. Barbagallo1, R. Barlow25, C. Bartoli26, N. Bartosik9, E. Barzi13, F. Batsch1, M. Bauce11, M. Begel27, J. S. Berg27, A. Bersani28, A. Bertarelli1, F. Bertinelli1, A. Bertolin10, P. Bhat13, C. Bianchi26, M. Bianco1, W. Bishop17,29, K. Black30, F. Boattini1, A. Bogacz31, M. Bonesini32, B. Bordini1, P. Borges de Sousa1, S. Bottaro33, L. Bottura1, S. Boyd17, M. Breschi26,6, F. Broggi22, M. Brunoldi34,4, X. Buffat1, L. Buonincontri35,10, P. N. Burrows36, G. C. Burt37,38, D. Buttazzo39, B. Caiffi28, S. Calatroni1, M. Calviani1, S. Calzaferri34, D. Calzolari1, C. Cantone40, R. Capdevilla13, C. Carli1, C. Carrelli41, F. Casaburo11,42, M. Casarsa43, L. Castelli42,11, M. G. Catanesi44, L. Cavallucci26,6, G. Cavoto42,11, F. G. Celiberto45, L. Celona46, A. Cemmi41, S. Ceravolo40, A. Cerri47,48,39, F. Cerutti1, G. Cesarini40, C. Cesarotti49, A. Chancé23, N. Charitonidis1, M. Chiesa4, P. Chiggiato1, V. L. Ciccarella40,42, P. Cioli Puviani50, A. Colaleo51,44, F. Colao41, F. Collamati11, M. Costa52, N. Craig53, D. Curtin54, L. D’Angelo55, G. Da Molin56, H. Damerau1, S. Dasu30, J. de Blas57, S. De Curtis58, H. De Gersem55, T. Del Moro42,41, J.-P. Delahaye1, D. Denisov27, H. Denizli59, R. Dermisek60, P. Desiré Valdor1, C. Desponds1, L. Di Luzio10, E. Di Meco40, K. F. Di Petrillo18, I. Di Sarcina41, E. Diociaiuti40, T. Dorigo10,61, K. Dreimanis62, T. du Pree63,64, T. Edgecock25, S. Fabbri1, M. Fabbrichesi43, S. Farinon28, G. Ferrand23, J. A. Ferreira Somoza1, M. Fieg65, F. Filthaut66,63, P. Fox13, R. Franceschini67,68, R. Franqueira Ximenes1, M. Gallinaro56, M. Garcia-Sciveres3, L. Garcia-Tabares69, R. Gargiulo42, C. Garion1, M. V. Garzelli70, M. Gast71, C. E. Gerber72, L. Giambastiani35,10, A. Gianelle10, E. Gianfelice-Wendt13, S. Gibson7, S. Gilardoni1, D. A. Giove22, V. Giovinco1,C. Giraldin35,10, A. Glioti23, A. Gorzawski73,1, M. Greco68, C. Grojean74, A. Grudiev1, E. Gschwendtner1, E. Gueli11,11, N. Guilhaudin1, C. Han75, T. Han76, J. M. Hauptman20, M. Herndon30, A. D. Hillier29, M. Hillman77, T. R. Holmes77, S. Homiller78, S. Jana79, S. Jindariani13, S. Johannesson73, B. Johnson77, O. R. Jones1, P.-B. Jurj80, Y. Kahn13, R. Kamath80, A. Kario64, I. Karpov1, D. Kelliher29, W. Kilian81, R. Kitano82, F. Kling74, A. Kolehmainen1, K. C. Kong83, J. Kosse16, G. Krintiras83, K. Krizka84, N. Kumar85, E. Kvikne1, R. Kyle86, E. Laface73, K. Lane87, A. Latina1, A. Lechner1, J. Lee19, L. Lee77, S. W. Lee19, T. Lefevre1, E. Leonardi11, G. Lerner1, P. Li88, Q. Li89, T. Li90, W. Li91, R. Li Voti42,40, M. Lindroos73, R. Lipton13, D. Liu76, M. Liu1, Z. Liu88, A. Lombardi1, S. Lomte30, K. Long80,29, L. Longo44, J. Lorenzo92, R. Losito1, I. Low93,94, X. Lu17, D. Lucchesi35,10, T. Luo3, A. Lupato35,10, E. Métral1, K. Mękała95,74, Y. Ma6, J. M. Mańczak1, S. Machida29, T. Madlener74, L. Magaletti96,44, M. Maggi44, H. Mainaud Durand1, F. Maltoni97,26,6, M. Mandurrino9, C. Marchand23, F. Mariani22,42, S. Marin1, S. Mariotto21,22, S. Martin-Haugh29, M. R. Masullo98, G. S. Mauro46, A. Mazzolari24,99, B. Mele11, F. Meloni74, X. Meng100, M. Mentink1, R. Miceli26, N. Milas73, A. Mohammadi30, D. Moll55, A. Montella101, M. Morandin10, M. Morrone1, T. Mulder1, R. Musenich28, M. Nardecchia42,43, F. Nardi35, D. Neuffer13, D. Newbold29, D. Novelli28,42, M. Olvegård102, Y. Onel103, D. Orestano67,68, J. Osborne1, S. Otten64, Y. M. Oviedo Torres104, D. Paesani40,1, S. Pagan Griso3, D. Pagani6, K. Pal1, M. Palmer27, A. Pampaloni28, P. Panci39,105, P. Pani74, Y. Papaphilippou1, R. Paparella22, P. Paradisi35,10, A. Passeri68, N. Pastrone9, A. Pellecchia44, F. Piccinini4, H. Piekarz13, T. Pieloni106, J. Plouin23, A. Portone92, K. Potamianos17, J. Potdevin106,1, S. Prestemon3, T. Puig107, J. Qiang3, L. Quettier23, T. R. Rabemananjara108,63, E. Radicioni44, R. Radogna51,44, I. C. Rago11, A. Ratkus62, E. Resseguie3, J. Reuter74, P. L. Ribani26, C. Riccardi34,4, S. Ricciardi29, T. Robens109, Y. Robert1, C. Rogers29, J. Rojo63,108, M. Romagnoni99,24, K. Ronald86,38, B. Rosser18, C. Rossi1, L. Rossi21,22, L. Rozanov18, M. Ruhdorfer110, R. Ruiz111, F. S. Queiroz104,111, S. Saini47,1, F. Sala26,6, C. Salierno26, T. Salmi112, P. Salvini4,34, E. Salvioni47, N. Sammut113, C. Santini22, A. Saputi24, I. Sarra40, G. Scarantino22,42, H. Schneider-Muntau114, D. Schulte1, J. Scifo41, T. Sen13, C. Senatore115, A. Senol59, D. Sertore22, L. Sestini10, R. C. Silva Rêgo104,111, F. M. Simone96,44, K. Skoufaris1, G. Sorbello116,46, M. Sorbi21,22, S. Sorti21,22, L. Soubirou23, D. Spataro74, A. Stamerra51,44, S. Stapnes1, G. Stark117, M. Statera22, B. M. Stechauner118,1, S. Su119, W. Su75, X. Sun89, A. Sytov24, J. Tang75, J. Tang120,100, R. Taylor1, H. Ten Kate64,1, P. Testoni92, L. S. Thiele2,1, R. Tomas Garcia1, M. Topp-Mugglestone1,36, T. Torims62,1, R. Torre28, L. T. Tortora68,67, S. Trifinopoulos49, S.-A. Udongwo2,1, I. Vai34,4, R. U. Valente22, U. van Rienen2, R. van Weelderen1, M. Vanwelde1, G. Velev13, R. Venditti51,44, A. Vendrasco77, A. Verna41, A. Verweij1, P. Verwilligen44, Y. Villamzar104, L. Vittorio121, P. Vitulo34,4, I. Vojskovic73, D. Wang89, L.-T. Wang18, X. Wang122, M. Wendt1, M. Widorski1, M. Wozniak1, Y. Wu123, A. Wulzer124,125, K. Xie76, Y. Yang126, Y. C. Yap74, K. Yonehara13, H. D. Yoo127, Z. You75, M. Zanetti35, A. Zaza51,44, L. Zhang86, R. Zhu128,129, A. Zlobin13, D. Zuliani35,10, J. F. Zurita130
1 European Organization for Nuclear Research, Geneva, Switzerland,
2 University of Rostock, Rostock, Germany,
3 Lawrence Berkeley National Laboratory, Berkeley, CA, United States,
4 INFN Sezione di Pavia, Pavia, Italy,
5 Ankara University, Ankara, Türkiye,
6 INFN Sezione di Bologna, Bologna, Italy,
7 Royal Holloway University of London, Egham, United Kingdom,
8 University of Turin, Turin, Italy,
9 INFN Sezione di Torino, Turin, Italy,
10 INFN Sezione di Padova, Padua, Italy,
11 INFN Sezione di Roma I, Rome, Italy,
12 University of Manchester, Manchester, United Kingdom,
13 Fermi National Accelerator Laboratory, Batavia, IL, United States,
14 University of Oregon, Eugene, OR, United States,
15 Fayoum University, Al Fayyūm,, Egypt,
16 Paul Scherrer Institute, Villigen, Switzerland,
17 University of Warwick, Coventry, United Kingdom,
18 University of Chicago, Chicago IL, United States America,
19 Kyungpook National University, Daegu, South Korea,
20 Iowa State University, Ames, IA, United States,
21 University of Milan, Milan, Italy,
22 INFN Laboratorio LASA, Segrate, Italy,
23 Institut de recherche sur les lois fondamentales de l’Univers, Gif-sur-Yvette, France,
24 INFN Sezione di Ferrara, Ferrara ,Italy,
25 University of Huddersfield, Huddersfield, United Kingdom,
26 University of Bologna, Bologna, Italy,
27 Brookhaven National Laboratory, Upton, NY, United States,
28 INFN sezione di Genova, Genova, Italy,
29 ISIS Neutron and Muon Source, Didcot, United Kingdom,
30 University of Wisconsin-Madison, Madison, WI, United States,
31 Thomas Jefferson National Accelerator Facility, Newport News, VA, United States,
32 INFN Sezione di Milano, Milan,Italy,
33 Tel Aviv University, Tel Aviv, Israel,
34 University of Pavia, Pavia, Italy,
35 University of Padua, Padua, Italy,
36 John Adams Institute for Accelerator Science, Oxford, United Kingdom,
37 Lancaster University, Lancaster, United Kingdom,
38 Cockcroft Institute, Daresbury, United Kingdom,
39 INFN Sezione di Pisa, Pisa, Italy,
40 INFN National Laboratory of Frascati, Frascati, Italy,
41 National Agency for New Technologies, Energy and Sustainable Economic Development, Rome, Italy,
42 Sapienza University of Rome, Rome,Italy,
43 INFN Sezione di Trieste, Trieste, Italy,
44 INFN Sezione di Bari, Bari, Italy,
45 Universidad de Alcalá, Alcalá de Henares, Spain,
46 INFN Laboratori Nazionali del Sud, Catania, Italy,
47 University of Sussex, Brighton, United Kingdom,
48 University of Siena, Siena, Italy,
49 Massachusetts Institute of Technology, Cambridge, MA, United States,
50 Polytechnic University of Turin, Turin, Italy,
51 University of Bari, Bari, Italy,
52 Perimeter Institute, Waterloo, ON, Canada,
53 University of California, Santa Barbara, Santa Barbara, CA, United States,
54 University of Toronto, Toronto, ON, Canada,
55 Technical University of Darmstadt, Darmstadt, Germany,
56 LIP—Laboratory of Instrumentation and Experimental Particle Physics, Lisbon, Portugal,
57 Universidad de Granada, Granada, Spain,
58 INFN Sezione di Firenze, Sesto Fiorentino, Italy,
59 Bolu Abant Izzet Baysal University, Bolu, Türkiye,
60 Indiana University, Bloomington, IN, United States,
61 Luleȧ University of Technology, Luleȧ, Sweden,
62 Riga Technical University, Riga, Latvia,
63 National Institute for Subatomic Physics, Amsterdam, Netherlands,
64 University of Twente, Enschede, Netherlands,
65 University of California, Irvine, Irvine, CA, United States,
66 Radboud University Nijmegen, Nijmegen, Netherlands,
67 Roma Tre University, Rome, Italy,
68 INFN Sezione di Roma III, Rome, Italy,
69 Centro de Investigaciones Energéticas, Medioambientales y Tecnologicas, Madrid, Spain,
70 Universität Hamburg, Hamburg, Germany,
71 Karlsruhe Institute of Technology, Karlsruhe, Germany,
72 University of Illinois at Chicago, Chicago, IL, United States,
73 European Spallation Source, Lund, Sweden,
74 Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany,
75 Sun Yat-sen University, Guangzhou, China,
76 University of Pittsburgh, Pittsburgh, PA, United States,
77 University of Tennessee at Knoxville, Knoxville, TN, United States,
78 Harvard University, Cambridge, MA, United States,
79 Max Planck Institute for Nuclear Physics, Heidelberg, Germany,
80 Imperial College London, London, United Kingdom,
81 University of Siegen, Siegen, Germany,
82 High Energy Accelerator Research Organization, Tsukuba, Japan,
83 University of Kansas, Lawrence, KS, United States,
84 University of Birmingham, Birmingham, United Kingdom,
85 Shree Guru Gobind Singh Shree Guru Gobind Singh Tricentenary University, Gurgaon, India,
86 University of Strathclyde, Glasgow, United Kingdom,
87 Boston University, Boston, MA, United States,
88 University of Minnesota, Minneapolis, MN, United States,
89 Peking University, Beijing, China,
90 Nankai University, Tianjin, China,
91 Rice University, Houston, TX, United States,
92 Fusion for Energy, Barcelona, Spain,
93 Northwestern University, Evanston, IL, United States,
94 Argonne National Laboratory, Lemont, IL, United States,
95 University of Warsaw, Warsaw, Poland,
96 Polytechnic University of Bari, Bari, Italy,
97 UCLouvain, Louvain-la-Neuve, Belgium,
98 INFN Sezione di Napoli, Naples, Italy,
99 University of Ferrara, Ferrara, Italy,
100 Institute of High Energy Physics, Beijing, China,
101 Stockholm University,Stockholm, Sweden,
102 Uppsala University, Uppsala, Sweden,
103 University of Iowa, Iowa City, IA, United States,
104 Universidade Federal do Rio Grande do Norte, Campus Universitário, Natal, Brazil,
105 University of Pisa, Pisa, Italy,
106 École polytechnique fédérale de Lausanne, Lausanne, Switzerland,
107 Institut de Ciencia de Materials de Barcelona, Cerdanyola del Vallès, Spain,
108 Vrije Universiteit Amsterdam, Amsterdam, Netherlands,
109 Rudjer Boskovic Institute, Zagreb, Croatia,
110 Cornell University, Ithaca, NY, United States,
111 Institute of Nuclear Physics, Polish Academy of Sciences, Kraków, Poland,
112 Tampere University, Tampere, Finland,
113 University of Malta, Msida, Malta,
114 CS&T, France,
115 University of Geneva, Geneva, Switzerland,
116 University of Catania, Catania, Italy,
117 University of California, Santa Cruz, Santa Cruz, CA, United States,
118 TU Wien, Vienna, Austria,
119 University of Arizona, Tucson, AZ, United States,
120 University of Science and Technology of China, Hefei, China,
121 Laboratoire d’Annecy de Physique des Particules LAPP, Annecy, France,
122 University of California, San Diego, La Jolla, CA, United States,
123 Nanjing Normal University, Nanjing, China,
124 Institució Catalana de Recerca i Estudis Avançats, Barcelona, Spain,
125 Institut de Física d’Altes Energies (IFAE), Barcelona, Spain,
126 University of Southampton, Southampton, United Kingdom,
127 Yonsei University, Seoul, South Korea,
128 Institute of Modern Physics, Lanzhou, China,
129 University of Chinese Academy of Sciences, Beijing, China,
130 Instituto de Física Corpuscular, Paterna, Spain
Executive summary
The recent European Strategy for Particle Physics Update (ESPPU) and the Snowmass processes have firmly established the need to get to the 10 TeV partonic center-of-mass (pCOM) energy collisions in order to make a major step forward in our understanding of particle physics. The International Muon Collider Collaboration (IMCC) was formed in 2020 to secure a novel, sustainable path toward this goal by developing the muon collider concept and addressing the related challenges. The collaboration obtained strong support by many partner institutes and funding agencies across the globe, the European Union and recently the Particle Physics Project Prioritization Panel (P5) in the US. It is currently hosted at CERN and has 50 full and several associated members. IMCC aims to develop a staged approach to a 10 TeV, high-luminosity muon collider.
The report summarizes progress made since the inception of IMCC and describes what is expected to be achieved by 2026 with currently secured resources. It also outlines important additional accelerator and detector studies that could be addressed and completed by 2026, if additional resources were made available. In 2025, IMCC will provide the ESPPU with an evaluation report that assesses the muon collider potential and an R&D path plan. Later it will submit the updated information to strategy process in the U.S. and potentially other regions and countries.
Motivation
A muon collider with 10 TeV energy would enable an exceptionally broad physics programme. It could discover new particles with presently inaccessible mass, including weakly interacting massive particles (WIMP) dark matter candidates. It could discover cracks in the Standard Model by the precise study of the Higgs boson, including the direct observation of double-Higgs production and the precise measurement of triple Higgs coupling. It will uniquely pursue the quantum imprint of new phenomena in novel observables by combining precision with energy. It gives unique access to new physics coupled to muons and delivers beams of neutrinos with unprecedented properties from the muons’ decay.
This unmatched physics program can be realised because the muon collider offers major advantages in terms of the footprint, cost, and power consumption when compared to the next generation of high-energy lepton and hadron machines. In order to directly probe the existence of new heavy states, the luminosity of colliders has to increase with the square of the collision energy to compensate for the reduction in $s$-channel cross sections. The potential of muon colliders to improve the luminosity to beam power ratio at high energies is therefore a major benefit. Based on a combination of physics and technical considerations, the integrated luminosity target was set to 10 ab-1 at 10 TeV.
Work programme and resources
The IMCC, the muon beam panel of the European Large National Laboratories Directors Group (LDG) and the Snowmass process in the U.S. have all assessed the muon collider concept. They concluded that the concept is less mature than linear colliders but that there are no insurmountable obstacles. The concept promises a unique path toward high energy and high luminosity with limited cost and power consumption; this makes it important to address the challenges of the concept and open this path to the future.
A vigorous R&D program is needed to develop an end-to-end design and to unequivocally demonstrate feasibility of the machine and the detector performance. With involvement of the global community, a concise set of work-packages has been developed for the European Strategy for Particle Physics—Accelerator R&D Roadmap by the Laboratory Directors Group, hereinafter referred to as the European LDG roadmap. This process provided an excellent basis for the global collaboration and planning of future work. The short-term goal is to assess initial technical feasibility of a 10 TeV collider, identify key R&D items, develop an implementation timeline, and provide an initial cost estimate in time for the next ESPPU process. Further studies will address strategic needs in the U.S. and other regions.
IMCC is actively pursuing resources to complement the resources already foreseen by CERN and from other collaboration partners. In 2023 European Commission support was obtained for a design study of a muon collider. This project started in March 2023 and its work-packages are well aligned with the overall IMCC studies. The secured resources are used following the priorities established in the European LDG roadmap and will be able to cover about 40% of the programme by the end of 2026.
The project continues to make important progress and the collaboration steadily grows as detailed in this report. Examples of the ongoing efforts include studies and prototyping plans of the magnet and RF systems and their integration into a muon cooling cell. Progress is being made in designs of the muon production target, the muon cooling section, the fast-ramping accelerators, and the collider ring. Detailed studies of the machine-detector interface and the beam-induced background in the detector are also advancing.
A further increase in the design effort for the accelerator complex and the detector as well as the associated technologies is required to achieve the goals presented in the European LDG roadmap. In addition, the start of an experimental program is essential to develop the technologies, in particular the muon cooling cell magnets, RF systems and absorbers as well as their integration. Other key items are the muon production target and its solenoid, which is similar to what is required for fusion reactors. The fast-ramping magnets and power converter of the acceleration system. Several test infrastructures are instrumental to achieve progress, including one to test RF cavities in high magnetic field and one to test a muon cooling module prototype.
In December 2023, the U.S. Particle Physics Projects Prioritization Panel (P5) recommended that the U.S. should develop a collider with 10 TeV parton collision energies, such as a muon collider, a proton collider, or possibly an electron-positron collider. The report states: “In particular, a muon collider presents an attractive option both for technological innovation and bringing the energy frontier back to the U.S. The U.S. should participate to the IMCC and have the ambition to host a future collider.” Several U.S. institutes have already officially joined the IMCC and members of the U.S. community are actively engaged in the collaboration’s work. The amount of dedicated resources in the U.S. is not finalized and a considerable amount of time will be needed to ramp up the effort. IMCC plans to revise both the organization and the work plan together with colleagues from the U.S. and other regions. The ambition of the U.S. to host a muon collider facility further strengthens motivation for the R&D and facilitates the demonstrator planning at a faster pace. Besides the U.S., collaborative ties are being established with other regions of the world.
Site and timeline
The study is currently focusing on a site-agnostic design. The benefits of reusing the existing infrastructure, such as the existing proton complexes and collider tunnels, will be considered but detailed studies will only follow at a later time. Potential sites in Europe near CERN and in the United States at Fermilab are being explored and other regions may propose alternative sites as the study progresses. Close to CERN, a potential location and orientation of the collider ring have been identified that can mitigate effects of the neutrino flux. Muon collider implementation at Fermilab requires that the accelerator rings all fit within the laboratory’s boundaries and first considerations indicate that this is possible.
There are many global uncertainties in planning of the next generation of collider facilities. These include budgetary and geopolitical considerations that are often outside of control for our field. The muon collider is an important concept in the long run since it allows highest lepton collision energies; it could operate after or in parallel to a Higgs factory. In the U.S., the muon collider can be the next high-energy frontier project for particle physics. If no Higgs factory is being built Europe, the muon collider could be the next project after the High-Luminosity Large Hadron Collider (HL-LHC). In light of these uncertainties, we focus on development of the fastest implementation timeline, while maintaining as much flexibility as possible. This timeline is based on technical considerations, such as the typical amount of time needed to make one iteration on the magnet technology, and hence we refer to it as “technically limited”. We assume that no show-stoppers will occour during the course of the R&D programme and that the required funding and resources will be made available.
We consider a staged implementation of the collider, based on the anticipated maturity of the relevant technologies in about 15 years, in particular the most critical ones: the muon cooling technology, the detector technologies and the high temperature superconductor (HTS) solenoid technology. Only HTS-based collider ring magnet technology may not be mature at this timescale and is not planned for the first stage. Depending on the physics needs and the funding situation, energy staging or luminosity staging can be considered. One can start with a cheaper, lower energy stage (e.g. 3 TeV) at full luminosity performance that is later extended to 10 TeV or higher with improved technology. Alternatively, one can implement a 10 TeV collider directly, albeit with reduced luminosity, which can later be upgraded to the nominal luminosity, similarly to HL-LHC. Both approaches do not require HTS dipoles and allow a fast implementation of a muon collider with a start of operation before 2050, provided the decision-making process is well prepared and development resources are made available. A muon collider can thus be the next high-energy frontier flagship project in all regions, e.g. in Europe it could follow directly after the HL-LHC.
Synergies and outreach
Many young scientists have developed their scientific and technical skills in the field of particle physics and the associated accelerators; they also learned to work in large international collaborations. The muon collider is a novel concept and opens opportunities for junior researches to make innovative contributions that are much harder to make in long-established design approaches.
The muon collider has synergies with other particle and nuclear physics projects, e.g. in the detector technologies and concepts, the magnet technology for hadron colliders, the high-efficiency RF technology and high-power targets for neutron spallation sources. It also needs several technologies that differ from other colliders. High-field solenoids based on high-temperature superconductor are a prime example and are also of interest for fusion reactors and power generators for off-shore windmills as well as life and material sciences, e.g. nuclear magnetic resonance and magnetic reasonance imaging.
The test facility and the collider itself require a high power proton source. This allows to share technology and potentially even facilities with neutrino facilities (e.g. NuSTORM, neutrino factory, DUNE), lepton flavour violation experiments (e.g. Mu2e, Mu2e-II, COMET, AMF) and the next generation of low-energy, highly polarized muon beams.
Conclusion
The unparalleled physics case coupled with sustainability arguments created a very strong recent interest in the muon collider concept. This interest led to formation of the International Muon Collider Collaboration. The near term goal for the collaboration is to assess initial technical feasibility of the machine, identify key R&D items, develop an implementation timeline, and provide an initial cost estimate in time for the next ESPPU process. There has been a significant R&D progress made by the collaboration and its partner institutions, including those from the U.S. during the Snowmass process. The recent P5 recommendations further stressed the importance of the R&D program. Realization of a staged muon collider appears promising but the effort is in strong need for additional resources, in particular to strengthen the hardware developments.
0.1 Overview of collaboration goals, challenges and R&D programme
The International Muon Collider Collaboration (IMCC) [1] was established in 2020 following the recommendations of the European Strategy for Particle Physics (ESPP) and the implementation of the European Strategy for Particle Physics—Accelerator R&D Roadmap by the Laboratory Directors Group [2], hereinafter referred to as the the European LDG roadmap. The Muon Collider Study (MuC) covers the accelerator complex, detectors and physics for a future muon collider. In 2023, European Commission support was obtained for a design study of a muon collider (MuCol) [3]. This project started on 1st March 2023, with work-packages aligned with the overall muon collider studies. In preparation of and during the 2021–22 U.S. Snowmass process, the muon collider project parameters, technical studies and physics performance studies were performed and presented in great detail. Recently, the P5 panel [4] in the U.S. recommended a muon collider R&D, proposed to join the IMCC and envisages that the U.S. should prepare to host a muon collider, calling this their “muon shot”. In the past the U.S. Muon Accelerator Programme (MAP) [5] has been instrumental in studies of concepts and technologies for a muon collider.
0.1.1 Motivation
High-energy lepton colliders combine cutting edge discovery potential with precision measurements. Because leptons are point-like particles in contrast to protons, they can achieve comparable physics at lower centre-of-mass energies [6, 7, 8, 9]. However, to efficiently reach the 10+ TeV scale recognized by ESPP and P5 as a necessary target requires a muon collider. A muon collider with 10 TeV energy or more could discover new particles with presently inaccessible mass, including WIMP dark matter candidates. It could discover cracks in the Standard Model (SM) by the precise study of the Higgs boson, including the direct observation of double-Higgs production and the precise measurement of triple Higgs coupling. It will uniquely pursue the quantum imprint of new phenomena in novel observables by combining precision with energy. It gives unique access to new physics coupled to muons and delivers beams of neutrinos with unprecedented properties from the muons’ decay. Based on physics considerations, an integrated luminosity target of 10 ab-1 at 10 TeV was chosen. However, various staging options are possible that allow fast implementation of a muon collider with a reduced collision energy or the luminosity in the first stage and reaches the full performance in the second stage.
In terms of footprint, costs and power consumption a muon collider has potentially very favourable properties.
Refer to captionRefer to caption
Figure 0.1.1: Left: Conceptual scheme of the muon collider. Right: Comparison of CLIC and a muon collider luminosities normalised to the beam power and as a function of the centre-of-mass energy.
The luminosity of lepton colliders has to increase with the square of the collision energy to compensate for the reduction in $s$-channel cross sections. Figure 0.1.1 (right panel) compares the luminosities of the Compact Linear Collider (CLIC) and a muon collider, based on the U.S. Muon Accelerator Programme (MAP) parameters [7], as a function of centre-of-mass energy. The luminosities are normalised to the beam power. The potential of muon colliders to improve the luminosity to beam power ratio at high energies is one of the main advantages of the concept.
0.1.2 The accelerator concept
IMCC studies a muon collider concept that has initially developed by MAP; a schematic view is shown in Fig. 0.1.1 (left panel).
The proton complex produces a short, high-intensity proton pulse that hits the target and produces pions. The decay channel guides the pions and collects the produced muons into a buncher and phase rotator system to form a muon beam. Several cooling stages then reduce the longitudinal and transverse emittance of the beam using a sequence of absorbers and RF cavities in a high magnetic field. A system of a linac and two recirculating linacs accelerate the beams to 63 GeV followed by a sequence of high-energy accelerator rings; the optimum sequence needs to be determined based on the ongoing studies. Finally the beams are injected at full energy into the collider ring. Here, they will circulate to produce luminosity until they are decayed; alternatively they can be extracted once the beam current is strongly reduced.
A set of parameters has been defined for 10 TeV and also 3 TeV. These are target parameters to explore the limits of each technology and design. If they can be fully met, the integrated luminosity goal could be reached within five years (or $2.5$ years, with two detectors) of full luminosity operation. This provides margin for further design and technology studies and a realistic ramp-up of the luminosity. It also enables to consider initial stages that can be implemented faster but often with reduced luminosity performance in this stage.
Table 0.1.1: Tentative target parameters for a muon collider at different energies. These values are only to give a first indication and correspond to the two stagin scenarios discussed in Section 0.1.6. The estimated luminosity refers to the value that can be reached if all target specifications can be reached; it also includes the beam-beam effect.
Parameter Symbol unit Scenario 1 Scenario 2
Stage 1 Stage 2 Stage 1 Stage 2
Centre-of-mass energy $E_{\mathrm{cm}}$ TeV 3 10 10 10
Target integrated luminosity $\int{\cal L}_{\mathrm{target}}$ $\rm ab^{-1}$ 1 10 10
Estimated luminosity ${\cal L}_{\mathrm{estimated}}$ $10^{34}\rm cm^{-2}s^{-1}$ 2.1 21 tbc 14
Collider circumference $C_{\mathrm{coll}}$ $\rm km$ 4.5 10 15 15
Collider arc peak field $B_{\mathrm{arc}}$ $\rm T$ 11 16 11 11
Luminosity lifetime $N_{\mathrm{turn}}$ turns 1039 1558 1040 1040
Muons/bunch $N$ $10^{12}$ 2.2 1.8 1.8 1.8
Repetition rate $f_{\mathrm{r}}$ $\rm Hz$ 5 5 5 5
Beam power $P_{\mathrm{coll}}$ $\rm MW$ 5.3 14.4 14.4 14.4
RMS longitudinal emittance $\varepsilon_{\parallel}$ $\rm eVs$ 0.025 0.025 0.025 0.025
Norm. RMS transverse emittance $\varepsilon_{\perp}$ µm 25 25 25 25
IP bunch length $\sigma_{z}$ $\rm mm$ 5 1.5 tbc 1.5
IP betafunction $\beta$ $\rm mm$ 5 1.5 tbc 1.5
IP beam size $\sigma$ µm 3 0.9 tbc 0.9
Protons on target/bunch $N_{\mathrm{p}}$ $10^{14}$ 5 5 5 5
Protons energy on target $E_{\mathrm{p}}$ $\rm GeV$ 5 5 5 5
BS photons $N_{\mathrm{BS},0}$ per muon 0.075 0.2 tbc 0.2
BS photon energy $E_{\mathrm{BS,0}}$ MeV 0.016 1.6 tbc 1.6
BS loss/lifetime (2 IP) $E_{\mathrm{BS,tot}}$ GeV 0.002 1.0 tbc 0.67
Table 0.1.2: Tentative target beam parameters along the acceleration chain. A 10 % emittance growth budget has been foreseen in the transverse and longitudinal planes, both for 3 and 10 TeV. This assumes that the technology and tuning procedures will have been improved between the two stages. The very first acceleration is assumed to be part of the final cooling. This choice allows to optimise the energy in the last absorber with no strong impact on the acceleration chain.
Parameter Symbol Unit Final cooling at 3 TeV at 10 TeV
Beam total energy $E_{\mathrm{beam}}$ GeV 0.255 1500 5000
Muons/bunch $N_{\mathrm{b}}$ $10^{12}$ 4 2.2 1.8
Longitudinal emittance $\varepsilon_{\parallel}$ $\rm eVs$ 0.0225 0.025 0.025
RMS bunch length $\sigma_{z}$ $\rm mm$ 375 5 1.5
[…]
Transverse norm. emittance $\varepsilon_{\perp}$ µm 22.5 25 25
[…]
0.1.3 Muon collider challenges
The Muon Collider Collaboration, the muon beam panel of the Laboratory Directors Group (LDG) and the Snowmass process in the U.S. have all assessed the muon collider challenges with the support of the global community. Key conclusions are that although the muon collider concept is less mature than several linear collider concepts no insurmountable obstacles have been identified, and that important design and technical challenges have to be addressed with a coherent international effort. Furthermore, past work, in particular within the U.S. Muon Accelerator Programme (MAP) [5], has demonstrated several key MuC technologies and concepts, and gives confidence that the overall concept is viable. Since then further component designs and technologies have been developed that provide increased confidence that one can cool the initially diffuse beam and accelerate it to multi-TeV energy on a time scale compatible with the muon lifetime. However, a fully integrated design has yet to be developed and further development and demonstrations of technology are required.
The current IMCC programme prepares the way towards a full conceptual design report (CDR) and a demonstration programme for a muon collider. The IMCC efforts cover physics potential as well as detector and accelerator design and performance studies. The programme will provide the necessary information and input for the next European Strategy for Particle Physics Update and other international strategy processes, allowing stakeholders to make informed decisions about these next steps.
Considering the facility design and technical challenges key performance drivers and goals have been identified, providing guidance for prioritising studies and efforts. They include the challenges that have previously identified by the MAP study but represent a wider set. Particularly important are:
  • Environmental impact. The compact footprint, limited cost and power consumption are intrinsic features that motivate the muon collider study in the first place. Radiation protection measures will ensure a negligible impact of the facility on the environment, similar to the LHC. Based on the experience of LHC and other accelerators, one can expect to mitigate the impact of losses in different parts of the accelerator complex. Particular attention will be paid to the neutrino flux that is produced by the decays of the muons in the collider ring and that exits the ground far from the collider. A mover system has been proposed that will avoid significant localised neutrino flux from the arcs by vertically changing the beam angle between $-1$ mradian to $1$ mradian and indeed can achieve a negligible level. The mover system is currently under study and the impact it has one the beam will be studied soon. The orientation of the The straight experimental insertions of the collider ring will lead to a localised higher neutrino flux. Civil engineering studies identified a position and an orientation of the collider ring in which the neutrinos will exit into the mediteranian sea and on the uninhabited side of a mountain, which could be fenced.
  • Machine-detector interface and detector. The muons of the beams that circulate in the collider ring decay, each producing two neutrinos and one electron or positron. The latter will hit the aperture and create showers. Tungsten masks protect the detector from this beam induced background. The detector will reduce the impact of the residual background by using components with high time and space resolution. Also the potential background from beam-beam effects is being explored but should be less severe. The studies indicate that the radiation in the detector is roughly similar to HL-LHC. A part of the physics measurements is not affected by the background, some part shows some residual impact. Further optimisation is ongoing to also remove this impact.
  • Proton complex. In the baseline, a proton beam power of around 2 MW at 5 Hz is used for muon production. Designs for proton facilities with similar or larger power exist. The main proton complex challenge arises from the combination of the protons into short, high-charge bunches. The corresponding workpackage of the collaboration is addressing this.
  • Muon production. The key challenge for the high-power target is the survival of the target itself under the shock waves of the incoming beam pulses and the temperature gradients to remove the deposited heat. Three technologies are under consideration, a solid graphite target, a liquid metal target or an fluidised tungsten target. The target is immersed into a 20 T solenoid field to efficiently collect the muons. This solenoid requires a large aperture to allow for sufficient shielding against the transverse showers from the target. Currently the studies indicate that a graphite target and its solenoid are feasible and can provide a sufficient number of muons.
    A liquid metal or fludised tungsten target would allow to sustain a higher proton beam power and produce a larger number of muons. This can lead to improved performance and provides margin with the muon bunch charge that could be used to reduce the cost of the downstream systems. The FCC-ee also plans to use a liquid metal target to dump the beamstrahlung photons from the interaction region.
  • Muon cooling design. Muon ionisation cooling increases the muon beam brightness by repeatedly slowing it in absorbers and re-accelerating it in RF cavities; both inside of strong solenoid fields to keep the beam focused. This principle has been demonstrated in MICE [10]. The target parameters anticipate some improvement of the cooling complex performance, e.g. a factor two reduction of the transverse emittance in the final cooling. Studies of the physical limitations of the cooling indicate that the target can be reached in principle. The current effort focuses on improving the lattice with the aim to reach this principle limit.
  • Muon cooling technology. The muon cooling system requires close integration of absorbers and RF cavities in a strong magnetic field. The operation of RF cavities in a magnetic field can strongly limit the gradient that can be achieved with no breakdown. This has been addressed by the MAP study, which showed that gradients in excess of the design target can be achieved by using cavities with beryllium endplates or by filling the cavities with hydrogen []. Unfortunately, the RF test stands in which these experiments were carried out has been dismanteled. The collaboration is developing a design for a new facility to test the RF in high magnetic field. IMCC is actively searching for the resources to implement such a facility and considers this a prime goals of the collaboration.
    The collaboration started to develop and engineering design of a cooling cell. This will address the challenges of its components and their integration into one unit. The findings may have important impact on the lattice design.
    The final cooling uses highest-field small aperture solenoids. Currently, fields of more than 30 T can be achieved using HTS in the whole solenoid. The goal for the next generation is 40 T and we use this value in the current design effort for the final cooling.
  • Muon acceleration. The lion share of the muon beam acceleration will be performed by a sequence of pulsed synchrotrons (RCS); an alternative use of fixed field accelerators (FFAs) is also considered. In each RCS the magnet field is ramped up in proportion to the energy gain of the beam. Some synchrotrons are based on a hybrid design where the fast-ramping magnets are interleaved with static superconducting ones. The pulsed synchrotrons face challenges in terms of optics design, the fast-ramping magnet systems and the RF systems. Field ramp rates between a few hundred T/s in the largest final and several kT/s in the smallest first ring are currently foreseen. The latter requires normal-conducting magnets while for the former superferric or HTS magnets can also be considered. Studies indicate that the ramp speeads can be obtained. A specific challenge is the large stored energy in the magnets (in total in the range of O(100 MJ)), which requires demanding power converters with very efficient recovery of the energy of each pulse for the subsequent one.
    Ongoing studies of the RCS considering lattice design, pulsed magnets and RF are promising. An integrated cost model is being developed that allows to identify the ramp shape with the best trade-off between RF and magnet demands.
  • Collider ring. The collider ring requires a small beta-function at the collision point, resulting in significant chromaticity that needs to be compensated. It also needs to maintain a short bunch.
    High-energy electrons and positrons that arise from muon decay and strike the collider ring magnets can cause radiation damage and unwanted heat load. Studies show that this can be mitigated with sufficient tungsten shielding and concepts of efficient cooling of these shields have been made. The shielding requires a larger magnet aperture in the range which is taken into account in their concept.
    A solution for 3 TeV collider ring lattice has been developed by the MAP study and successfully addresses the challenges. The design is based on Nb3Sn magnets with performances similar to the ones for HL-LHC. The magnet experts expect that this technology is fully mature in 15 years.
    A design of 10 TeV is more challenging because it requires an even smaller beta-function. The current studies with higher field HTS and hybrid magnets currently achieves the beta-function but does not yet achieve the target energy acceptance. Further efforts aim to improve this. Otherwise one would need to reduce the beam energy spread by about a factor two, potentially leading to a 30% luminosity reduction.
  • Collective effects. A very high muon bunch charge is required to achieve the luminosity goal. This can lead to important collective effects—such as space charge, beam loading, wakefields—that might limit the collider performance.
  • Cost and power consumption. The cost and power consumtion of the faciliy is an important challenge since sustainability must be prime concern for future scientific projects.
0.1.4 Developing the muon collider study
During and after the development of the European LDG roadmap, several steps were important to generate the framework for the muon collider studies:
  • The muon collider has been reviewed by the global community. Both, the muon beam panel of the European LDG roadmap as well as the Energy Frontier and Accelerator Frontier working groups of the U.S. Snowmass process used the world-wide expertise in assessing the challenges of the concept. This has been instrumental in identifying the required R&D programme.
  • The International Muon Collider Collaboration has been established. The collaboration provides the organisational framework for all the efforts to develop the collider concept.
  • The inclusion of the muon collider in the medium term plan of CERN in 2021, providing an important part of the resources.
  • The decision of the European Union to fund the proposal for a co-financed design study (MuCol) that started in 2023. This also served to initiate the additional contributions from many of the collaboration partners.
  • At the end of 2023, the U.S. P5 process recommended that the U.S. join the IMCC and to consider hosting such a facility. The collaboration organisation will adapt to fully include the U.S. as efforts will ramp up. The goals set out for the muon collider studies during the Snowmass process and in the P5 recommendations are being integrated into the IMCC future programme.
The European LDG roadmap identifies a prioritised R&D programme to make fully informed decisions at the next strategy process. It estimated the required resources to complete this programme to be a total of 446 FTE-years and 11.9 MEUR. With the currently secured funding the collaboration can provide slightly less than 200 FTE-years until and including 2025. Almost no material budget is available at this moment. However, the collaboration is activley seeking to increase the funding.
IMCC short-term goals
The International Muon Collider Collaboration has formed with the short term goal to address the European LDG roadmap in time for the next European Strategy for Particle Physics Update. More than 60 partner institutions are currently involved in the muon collider studies. Three main deliverables are foreseen:
  • a project evaluation report that assesses the muon collider potential;
  • a R&D plan that describes a path towards the collider;
  • an interim report early in 2024 (the present document) that documents progress and allows the wider community to be updated on the concept and to give feedback to the collaboration.
IMCC envisages to study the 10 TeV option, and also explore lower and higher energy options, e.g., a 3 TeV option as a step toward 10 TeV. The details of the required work and the required resources are documented in the European LDG roadmap [2].
Overall facility parameters as cost, power/energy consumption and carbon footprint will be estimated as the design matures, but are in many cases driving technical choices and optimisation.
IMCC organisation
A memorandum of cooperation (MoC) for the IMCC has been drawn up and the collaboration is rapidly growing. The formal structure of the collaboration, as shown in Fig. 0.1.2 consists of the following main bodies:
  • International collaboration board (ICB): The ICB role is to provide a forum for participants to examine ongoing activities, assure appropriate, well directed use of contributions and ensure a balanced portfolio of engagements. The chair of the ICB is elected by its members (simple majority). The ICB consists of one representative of each participating institute that has signed the MoC. The ICB has been active since October 2022.
  • Steering board (SB): The SB oversees the execution of the muon collider study by assessing the progress, resource usage and providing guidance. The SB reports to the LDG in Europe and to similar regional supervising organisations in other regions as required. The SG is active and operational since Spring 2023.
  • International advisory committee (IAC): The mandate of the IAC is to review the scientific and technical progress of the study. This body is active since Spring 2024.
  • Coordination committee (CC): The CC performs the overall coordination and execution of the muon collider study. It is chaired by the study leader. This body consists of the workpackage leaders and some additional experts of the collaboration and is very active. It provides the daily guidance of the study.
The workpackages of the collaboration are shown in Fig. 0.1.2.
Figure 0.1.2: Organigram of the muon collider study governance structure
The study leader is proposed by the host organization, and endorsed by the ICB. The study leader was appointed in 2022 and has led the work to build up the collaboration and activities since then.
The EU co-funded design study MuCol
IMCC successfully applied for an EU cofunded design study (named MuCol) [3]. The project started on $1^{\mathrm{st}}$ March 2023. The total co-funding of MuCol amounts to 3 MEUR, provided by the European Commission, the U.K. and Switzerland. In the design study, CERN only receives limited contributions for administrative support and travel, but has, in support of the successful design study bid, increased its contribution to the muon collider study.
The design study is fully integrated in the overall muon collaboration. The technical meetings and leaders are in common and governance and management are also synchronised. The organisation is shown in Fig. 0.1.3. Important activities in the collaboration are directly supported by the MuCol work-packages. In particular MuCol contains the workpackges Physics and Detector Requirements, Proton Complex, Muon Production and Cooling, High-energy Complex, Radio Frequency Systems, Magnet Systems and Cooling Cell Integration.
Figure 0.1.3: The MuCol work-packages and their interactions.
Extending the collaboration—US plans after P5
In the US, the Particle Physics Projects Prioritization Panel (P5) recommended in December 2023 that the U.S. should develop a collider with 10 TeV parton collision energies, such as a muon collider, a proton collider, or maybe an electron-positron collider. The report states: “In particular, a muon collider presents an attractive option both for technological innovation and bringing the energy frontier back to the US. The U.S. should participate to the IMCC and have the ambition to host a future collider.”
U.S. groups are very active and in leading roles in the IMCC, e.g., in the Coordination Committee and the Publications and Speakers Committee, even though in several cases the formal MoC signatures and effort ramp-up the will need to wait for the forthcoming implementation planning of the P5 recommendations.
The collaboration plans to revise both the organisation and the workplan together with U.S. colleagues. This will also impact the work programme laid out in this report. In particular, the ambition of the U.S. to host such a facility strengthens the motivation for R&D and demonstrator planning at a fast pace.
The collaboration also plans to attempt further strengthen other regions’ participation during this process. There is at this moment an important potential for substantial expansion of the collaboration, resources and efforts in Asia, and several European countries, Canada and South America.
0.1.5 R&D programme
The novelty of the muon collider concept implies that the R&D programme contains more uncertainty and challenges than for more conservative collider approaches.
A concise set of work-packages has been developed for the European LDG roadmap [2]. The programme aims to provide a broad basis for the global collaboration and planning of future work towards a multi-TeV muon collider. The programme is estimated to require about 450 full-time equivalent (FTE) years of labour and 12 MEUR material budget for the accelerator studies.
Substantial progress has been made since the definition of the European LDG roadmap. The approval of the EU co-funded design study MuCol, contributions from the collaboration members as well from U.S. partners during the Snowmass process and an increase of the budget at CERN were instrumental in supporting the muon collider study.
With the existing funding until 2026, the R&D cannot cover the full programme as described in the European LDG roadmap. The focus is on the most critical challenges and is guided by the priorities established by the roadmap. The planned effort is roughly consistent with the minimal scenario in the roadmap.
The focus is on the most critical challenges:
  • The development of the physics case.
  • The possibility to find a site and mitigate the neutrino flux.
  • The detector concept, in particular, the machine-detector interface which drives the physics capabilities of the detector.
  • The development of the high energy accelerator design concepts, in particular the pulsed synchrotrons and the collider ring. These are vital to reach the high energy beyond the scale studied by MAP.
  • The further development of the muon production and cooling complex. The focus is on an engineering design of the cooling cell and the evaluation of the target concept. An overall optimisation of the complex is desirable but very resource demanding.
  • The magnet technologies that drive the performance, cost and power consumption.
  • The RF technologies that drive the performance, cost and power consumption.
  • Other key technologies, for example shielding and cryogenics.
Work organisation
The distribution of the work onto workpackages has been modified since the publication of the European LDG roadmap to adjust to the availability of effort and experts. The document reflects this change. The areas represented in the coordination committee are:
  • The physics potential links to all studies of the muon collider physics potential.
  • The detector and MDI Links to the detector design studies and to the machine detector interface.
  • The proton complex addresses the key challenge of the accumulation of the protons in very high-charge bunches, by addressing in detail the proton complex design, and provides the basic parameters of the complex and the characteristics of the beam impacting on the production target.
  • The muon production and cooling addresses the production of the muons by the proton beam hitting a target and the subsequent cooling, including some of the specific technologies, such as for the production target and the absorbers that reduce the beam phase space volume.
  • The muon acceleration complex (in MuCol part of the high-energy complex) studies the acceleration complex of the muons.
  • The muon collider ring and siting studies the design of the collider ring. Because the muon collider siting is foremost driven by considerations on the neutrino flux from the collider ring the package also contains these activities.
  • Magnet systems will establish a complete inventory of the necessary magnets to optimize and standardise the design, and address the most critical ones. In particular it focuses on the solenoids of the muon production and cooling, which are specific to the muon collider, and the fast-ramping magnet system, which have ambitious requirements on power flow and power efficiency and limits the energy reach of the collider.
  • Radio frequency (RF) systems addresses the muon cooling ensuring coherence in frequency choice and synchronization among the various stages. It contributes to sustainability studies by its work on high efficiency RF power sources.
  • The beam-matter interaction and target design links to the different shielding studies, in particular for the muon production target and the collider ring.
  • The collective effects area includes all beam dynamics studies across the complex.
  • The demonstrator area connects to all the activities regarding the muon cooling demonstration.
  • Cooling cell integration addresses the design of the muon cooling cell, which is a unique and novel design and that faces integration challenges.
Recent progress and studies until 2026
The studies are summarised below, largely following the European LDG roadmap workpackages, which are indicated in brackets.
  • For the site and neutrino flux mitigation (MC.SITE, MC.NF) the following studies are ongoing and planned by 2026:
    • Verification of the neutrino flux impact using FLUKA and radiation protection knowledge.
    • Exploration of a site and orientation of the collider ring to mitigate the neutrino flux, in particular from the experimental insertions. This is the most difficult challenge.
    • Assessment of the mover concept to mitigate neutrino flux from the arcs.
    • An assessment if the neutrino flux mitigation with movers impacts beam operation.
    A number of important points cannot be covered with the current resources. This includes an assessment of the neutrino flux mitigation for the other parts of the complex, the development of a concept to measure and verify the neutrino flux in operation and the alignment concept that allows to link the machine reliably to the surface.
  • The MDI studies (MC.MDI) are progressing:
    • The simulation tools for the background prediction have been developed.
    • A masking system has been studied and optimisation is planned.
    • A programme to simulate beam-beam induced background has been developed.
    • Different lattice designs for the interaction region have been compared for the beam-induced background and limited differences were found.
    The main effort to mitigate the background will have to be carried by the detector design.
  • The accelerator concept and beam dynamics (MC.ACC) studies cover the key activities:
    • The combination of the high-power proton beam into short intense pulses.
    • A design of the experimental insertion and the arc lattices of the collider ring for 10 TeV.
    • A design of the arc lattice of the pulses synchrotrons of the muon beam acceleration system.
    • An improved concept of the muon cooling system.
    • Assessment of key collective effects and their impact on the performance and specifications.
    Studies that are not covered are the concepts of other parts of the collider complex, e.g., the low energy muon acceleration or the proton complex as well as full start-to-end simulations. Also alternatives for the muon cooling and acceleration cannot be covered. Further the studies do not cover the 3 TeV design. We anticipate that a solution for 10 TeV is also valid at 3 TeV.
  • The high-field magnet (MC.HFM) studies are ongoing
    • The muon production target solenoid concept is being developed, based on HTS technology.
    • A database with the magnet performance specifications from the previous MAP design has been developed and the resulting challenges identified. A tool to systematically derive optimised magnet configurations that create similar fields for the beam is under development and will be used on subsequent lattice designs.
    • 40 T has been identified as a realistic goal for the final cooling magnets and is used in the lattice design effort. The magnet study addresses the technology challenges that stem from this field.
    • Initial limits for the collider ring magnets have been identified and show the need for improved magnet protection and stress management, which will be addressed in further studies. Additional resources are required to perform more cable tests and to build magnet models to verify the simulations and improve the technology.
    An experimental programme would be important to verify magnet performances and to develop the technology, but the material budget for this is currently very limited. Also measurement of the radiation hardness of the HTS magnets would be instrumental in establishing the performance limits.
  • The design of the power converters and fast-ramping normal magnets (MC.FR) for the RCSs is ongoing in close interaction with the RF design and the lattice design. An overall optimisation will be performed of magnets, power converter and RF systems. These systems have a very important impact on cost and additional funding would allow experimental verification of the performances. The development of HTS-based alternatives is not covered.
  • The RF system (MC.RF) studies are ongoing:
    • The acceleration of high-charge muon bunches to high energies is being developed using superconducting cavities. The longitudinal beam dynamics studies along the high-energy complex is progressing and shows that cavities developed for the International Linear Collider (ILC) might be applicable.
    • The synchronisation is being studied between the RF and the ramping of the magnets in the RCS together with its impact on the cost.
    • The design is progressing of normal conducting cavities for the muon cooling cell.
  • The muon production target (MC.TAR) studies focus on the ability of the target to withstand the beam power and the required shielding of the surrounding solenoid. Indications are that a graphite target can sustain about 2 MW. Higher power liquid metal target alternatives will be explored. The current resources do not allow a complete design of the target and study of the muon capture systems.
  • The engineering design of one muon cooling cell (MC.MOD) is ongoing and is instrumental to maintain the option of a timely implementation of a demonstrator facility. Currently the resources to actually build models of the components and ultimately a prototype of the cell are not secured.
    Concepts of an RF test stand are being developed and allow to apply for funding to overcome the current lack of infrastructure of this type.
  • The initial scope for the muon cooling demonstrator (MC.DEM) has been determined and high-level concepts of the facility are being developed to understand the footprint requirements. Initially two potential sites at CERN have been identified, further options exist at Fermilab and we will invite more options. More resources are needed to develop a full conceptual design of the facility.
    The estimate of the cost scale (MC.INT) has started and will use a mixture of high-level estimates with detailed estimates of key components.
The physics potential and detector R&D goals are not part of the European LDG roadmap.
The physics potential requires sustained study to fully develop the understanding of the different measurements that can be performed at the muon collider, in particular also in view of synergies with other physics instruments.
A set of physics benchmark processes is being developed that allows to assess the key detector concept capabilities with a limited number of studies. The aim is to use these benchmark processes to demonstrate that the detector performance is adequate and that the background is well controlled.
Designs of the detector at 3 and 10 TeV detectors is being developed with clear specifications for the technology performance. A systematic review is important of these performance requirements and an assessment which current technology developments will meet them on which timescale. Also an exploration of alternative technologies is essential and the derivation of a prioritised list of alternatives that should be further developed. Of particular concern is the forward region, which suffers most from the background, as well as the detection of small angle muons. Also some concept for the luminosity monitoring is essential.
A 3 TeV detector concept exists based on a slightly modified version of the CLIC detector concept. Performance studies with this concept show that the measurement resolution for some processes is not strongly impacted by the beam-induced background while important impact exists for other channels. A 10 TeV detector concept is under development.
A systematic performance study of the 10 TeV detector concept is required for the benchmark processes to establish that the physics potential can be realised. Therefore simulation of the detector is required without and with beam-induced background in order to establish the latter has no significant impact on the physics performance. The development of artificial intelligence-based methods will be important to reduce the background at the reconstruction and potentially the trigger level.
0.1.6 Timeline and staging
The potential of a muon collider to reach 10 TeV parton-parton collisions with high luminosity makes it a very exciting future opportunity both on a longer and shorter timescale.
A roadmap towards a muon collider must remain flexible, and is an important part of a diverse international research programme in particle physics. New results, from the HL-LHC or non-accelerator based experiments, and potentially also lower energy experiments, might change the desired parameters for a future muon collider. Societal changes and priorities might also influence resource availability and hence timelines for investments in basic science as needed for future colliders.
We already know that sustainability and environmental consideration needs to guide the design process much more than projects decades back. Staging scenarios, a clear focus on working with developing technologies (from HTS magnets to machine learning (ML) and artificial intelligence (AI) methods) with a wide impact and interest outside our field, synergetic studies and projects as described in next subsection, and a broad international collaboration, are all important to develop the muon collider as a scientifically and technically novel but robust future facility when it comes to implementation.
The current timeline considerations identified that three technical developments are defining the minimum time to implement a muon collider:
  • The muon production and cooling technology and its demonstration in a test facility. At this moment we expect that this development can be mature enough for a decision in 15 years provided sufficient funding is made available.
  • The magnet technologies, in particular HTS superconductor technology. At this moment, we expect that the different HTS solenoids for the muon production and cooling are available within 15 years; the same is expected for the fast-ramping magnets. For the collider ring, one can expect 11 T Nb3Sn magnets with an aperture of 16 cm to be mature. Higher performant HTS or hybrid collider ring magnets may take longer.
  • The detector and its technologies that impact the efficiency of background suppression and the quality of the measurements. At this moment, we expect this technology to be mature in 15 years.
Sufficient funding should enable to accelerate the R&D of the other technologies and designs as to not constrain the timeline..
A staged implementation that anticipates use of Nb3Sn collider ring magnets can provide a muon collider by 2050. The scenarios are possible, see also Table 0.1.1:
  • In the energy-staging approach, the inital stage is at lower energy, for example 3 TeV. There is an important physics case already at this energy. In this strategy, the cost of the initial stage is substantially lower than for the full project. This could accelerate the decision-making processes. The 3 TeV design is consistent with Nb3Sn magnets at 11 T. In the second stage the whole complex will be reused with the exception of the collider ring. An RCS to accelerate to full energy and a new collider ring will be added.
  • In the luminosity-staging approach, the initial stage is at the full energy but using less performant collider ring magnets. This leads to an important reduction of the luminosity. If 11 T Nb3Sn dipoles are used instead of 16 T HTS dipoles, the resulting increase of collider ring circumference reduces the luminosity by one third. In addition a further reduction arises from the interaction regions. A detailed study is required to quantify the loss, but a factor three reduction might be a good guess. In the luminosity upgrade, the interaction region magnets will be replaced with more performant ones, similar to the HL-LHC. However one most likely will not replace the other collider ring magnets. In this scenario almost the complete project cost is required in the first stage, which could have important implications on the timeline.
The choice between the staging options will depend on physics needs and funding availability. Also the progress in magnet development will play an important role. Faster progress, in particular of HTS for the collider ring magnets, combined with strong funding support will make an earlier start of the 10 TeV option more attractive.
0.1.7 Site considerations
Different sites for the muon collider will be explored in the longer run. However, at this moment the international study to a large extent focuses on the technical and design challenges with no specific site in consideration. Specific sites however offer the benefit of existing infrastructure, e.g., existing proton complexes and collider tunnels. We will consider these benefits later for the different potential sites. Specific sites also can impose important constraints. For example, for Fermilab an option is preferred where the accelerator rings all fit on the existing site. First considerations indicate that this is possible. For CERN the most important consideration is to ensure that the neutrino flux can be mitigated to a negligible level. A first site and orientation of the collider ring that can achieve this has been identified and requires further study. In this case the extension of the experimental insertions point towards a steep uninhabited slope of the Jura on the one side and to the Mediterranean sea on the other side.
0.1.8 Synergies and outreach
Particle physics and the associated accelerator development have made important contributions to society; both in training of young people and in developing technologies.
Many young people have developed their scientific and technical skills in the field; they also learned to work in fully international collaborations. Because the muon collider is a novel concept it opens opportunities for young researchers to make original contributions to the development that are much harder to make in long-established design approaches.
The muon collider needs technologies in several areas that differ from other colliders. High-field solenoids are a prime example. In the past low-temperature superconductors such as the very mature NbTi and still developing Nb3Sn were the technologies of choice for accelerators and most other applications. Now high-temperature superconductors (HTS) are becoming an important technology. In particular they are of interest for fusion reactors, that have similar requirements than the one for the muon collider target solenoid. Highly-efficient superconducting motors and power generators, e.g., for off-shore windmills, also have strong synergy. Other relevant areas are life and material sciences; in particular, applications for nuclear magnetic resonance (NMR) and magnetic resonance imaging (MRI). In addition synergy exists with magnets for neutron spectroscopy, physics detectors and magnets for other particle colliders, such as hadron colliders.
The muon production target is synergetic with neutron spallation sources targets, in particular the alternative liquid metal concept.
The muon collider RF power sources have synergy with other developments of high-efficiency klystrons and superconducting cavities. Some RF systems need to work in high magnetic fields, an issue that also exists in some fusion reactor designs.
The test facility and the collider itself require a high power proton source. This allows sharing technology and potentially even facilities. Neutron spallation sources such as SNS and ESS are major examples; other examples are neutrino facilities, such a NuSTORM, lepton flavour violation experiments, such as mu2e and COMET and the generation of low-energy, highly polarised muon beams.